The Quick Buck Company is an all-equity firm that has been in existence for the past three years. Company management expects that the company will last for two more years and then be dissolved. The firm will generate cash flows of $900,000 next year and $1,400,000 in two years, including the proceeds from the liquidation. There are 35,000 shares of stock outstanding and shareholders require a return of 14 percent.

a. What is the current price per share of the stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Share price            $

b. The Board of Directors is dissatisfied with the current dividend policy and proposes that a dividend of $990,000 be paid next year. To raise the cash necessary for the increased dividend, the company will sell new shares of stock.

How many shares of stock must be sold? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Shares sold            

What is the new price per share of the existing shares of stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New share price            $
Answer #1

a.

Dividend per share in Year 1 = Cash flow generated next year / Number of shares

= $ 900,000 / 35,000

= $ 25.714

Dividend per share in Year 2 = Cash flow generated in two years / Number of shares

= $ 1,400,000 / 35,000

= $ 40

The share price today = [ Dividend in Year 1 / (1 + Required rate of return) ] + [ Dividend in Year 2 / (1 + Required rate of return)2 ]

= (25.714 / 1.14) + (40 / 1.142)

= $ 53.33

b.

Increase in the amount of dividend = $ 990,000 - $ 900,000 = $ 90,000

The number of new shares sold to pay the increase in the amount of dividend = Increase in dividend / Current price

= $ 90,000 / $ 53.33

= 1,687.6 shares

Dividend paid to the new shareholders in Year 2 = $ 90,000 * 1.14 = $ 102,600

Therefore, the amount of dividend paid to the existing shareholders in Year 2 will go down by $ 102,600

Dividend received by existing shareholders in Year 2 = 1,400,000 - 102,600 = $ 1,297,400

Dividend in Year 1 = $ 990,000 / 35,000 = 28.285

Dividend in Year 2 = $ 1,297,400 / 35,000 = 37.068

New price per share for existing shareholders

= (28.285 / 1.14) + (37.068 / 1.142)

= $ 53.33

The price is the same as calculated in part (a).